Instructions: Analyze each word problem. Write the equation you would use to solve the problem, but do not solve it.
The number of girls in Mrs. Conn's math class is 2 less than 3 times the number of boys. She has 26 students in her class. How many boys are in the class?
So basically, i'm trying to get the equation not the answer so can anyone help me please. i got 3n-2x=26???

Answer by KnightOwlTutor(293)

# of girls in class=3x-2
# of boys in class =x
We know that the number of students is 26
Therefore the equation will be 3x-2+x=26
simplify the x terms and add 2 to both sides
4x=28 divide both sides by 4 to get the # of boys

x=28/4 x=7
Although we are not to solve we can double chek our answer
7+3(7)-2= 5+21=26